# Quantum Enhanced Anomaly Detection for ADS-B Data using Hybrid Deep Learning

### v1

@@ I Introduction
Avionic systems have evolved significantly over the decades, transitioning from early federated architectures to today’s advanced Integrated Modular Avionics (IMA) frameworks. In the original federated model, each aircraft function—such as navigation, communication, or flight control—was managed by independent hardware and software components, leading to increased weight, complexity, and maintenance challenges. Modern IMA architectures have revolutionized this landscape by centralizing processing power through shared computing resources and modular software platforms, greatly enhancing flexibility, scalability, and fault tolerance [1]. This shift not only improves efficiency but also supports advanced communication protocols like Automatic Dependent Surveillance–Broadcast (ADS-B).
@@
In recent years, the adoption of ADS-B systems has led the aviation industry to an increase in the volume of aircraft surveillance data. The ADS-B system provides real-time data by transmitting the aircraft’ position, velocity and other flight characteristics to nearby aircraft and ground stations through a continuous, unencrypted data stream [2]. By leveraging satellite navigation, the ADS-B acquires precise positional information of an aircraft and is designed to improve aviation safety by enhancing situational awareness for pilots and air traffic controllers, aiding in collision avoidance and preventing runway incursions [3]. Furthermore, it helps optimize routing, improving the overall efficiency of air traffic management. However, since aviation systems receive broadcast ADS-B data without validating the source or integrity, the system is vulnerable to cyber threats such as spoofing, ghost aircraft injection, trajectory manipulation, and other attacks through the transmission of false data into air traffic control systems without validation [3],[4].
@@
Traditionally, a range of classical Machine Learning (ML) techniques have been employed in ADS-B security to detect anomalies and improve data validation. Algorithms such as Decision Trees, Support Vector Machines (SVM) and Neural Networks (NN) have demonstrated success in identifying and mitigating cyber threats in ADS-B data stream [5],[6]. However, these classical methods often involve high computational complexity and memory usage and may lose critical temporal information during feature transformation. These limitations highlight the need for more adaptive and scalable approaches to secure ADS-B communication, especially as air traffic surveillance data grows in complexity and volume.
@@
In parallel, as quantum hardware continues to improve, the growing field of QML has shown promising potential in accelerating the processing speed and effectively handling the high dimensionality associated with complex datasets through the quantum properties of superposition and entanglement [7],[8]. Hybrid quantum-classical models, which combine quantum circuits with classical neural networks, offer the potential to enhance learning capabilities while remaining compatible with near-term noisy intermediate-scale quantum (NISQ) devices.
@@
In this work, we explore the feasibility and application of a hybrid quantum-classical neural network, a H-FQNN, to perform anomaly detection on trajectory-based spoofing ADS-B data. We construct two quantum circuit-based hybrid models using one-hot encoded labels and binary cross-entropy loss and compare both to their classical counterpart, an FNN. The models are trained and evaluated on a publicly available ADS-B anomaly dataset that includes gradual manipulations of latitude and longitude, as well as merge attacks simulating new and complex flight patterns [5]. Detecting such trajectory anomalies is critical in operational aviation environments, as they may lead to false traffic alerts and in-flight disruptions, potentially resulting in serious safety consequences [9].
@@ II-A Quantum Computing (QC)
Quantum computing can be understood as a computational paradigm that leverages the principles of quantum mechanics and is built upon the fundamental unit of quantum information, the qubit [10]. Often represented as the spin of an electron or the polarization of a photon [11] [12], the qubit is the unit of computation with the basis states of:
@@
The fundamental utility of the quantum systems lies in the ability to represent information using the quantum mechanical property of entanglement and superposition. Superposition is the property that enables a quantum system to exist in multiple states simultaneously until measured. This can be represented as the linear combinations of states, written $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$ where $\alpha$ and $\beta$ are complex coefficients for which the square of their magnitudes, $|\alpha|^{2}$ and $|\beta|^{2}$, give the probabilities of measuring the system in the $|0\rangle$ and $|1\rangle$ states. Entanglement is another elementary property defining a group of qubits which are interconnected and cannot be described independently. In a system of 2 entangled qubits, it can be interpreted as the state of one qubit directly influencing the state of the other, with the famous example of the Bell state [13] [14]. A simple representation of the equation is:
@@ II-B Quantum Machine Learning (QML)
The convergence of Machine Learning and Quantum Computing has the potential to improve the communication technologies of the space and aeronautical sectors[15]. Quantum Machine Learning brings together the current state of the art in Machine Learning and harnesses the quantum mechanical properties of entanglement and superposition to attempt to create better performing models. There are four different approaches to combine QC and ML [16] [17] as illustrated in Figure 1. Classical input-Classical model (CC) refers to standard classical machine learning, where both data and models are classical. Classical input-Quantum model (CQ) involves processing classical data using quantum models, such as parameterized quantum circuits. Quantum input-Classical model (QC) describes the use of classical machine learning techniques to analyze data generated by quantum systems. Finally, Quantum input-Quantum model (QQ) leverages both quantum data and quantum models, enabling fully quantum learnable frameworks. These four paradigms define the various ways in which quantum and classical computing can synergize to advance machine learning capabilities. In this article, we refer to a hybrid model as a CQ framework that implements a quantum algorithm using classical input data.
Figure 1: Four different approaches to combine the disciplines of quantum computing and machine learning. (The figure is adapted from Schuld and Petruccione [16])
@@ II-C Variational Quantum Classifier (VQC)
A Variational Quantum Classifier can be considered as the quantum equivalent of a layer in a hybrid NN model. VQCs are particularly well suited for NISQ computers as the process of optimizing the cost function of a machine learning algorithm also helps mitigate the inherent decoherence and errors present in current quantum hardware [18]. The quantum circuit of a VQC begins with an initial state of qubits and uses a parameterized quantum circuit, known as a feature map, to encode features as quantum information [19] [20]. In this article, we used the angle embedding method, which encodes classical information of n features into the rotation angle of n qubits, effectively manipulating information from the vector space $\mathbb{C}^{N}$, where $N=2^{n}$ is the dimension of the Hilbert space corresponding to n qubits [21]. Specifically, we used $R_{X}$ gates, which apply a rotation of $\theta$ around the x-axis of a qubit, mapping the classical feature $x\in\mathbb{R}$ to the quantum state $|\psi(x)\rangle\in\mathcal{H}$, where $\mathcal{H}$ is a two-dimensional Hilbert space. The $R_{X}$ gate is defined by:
@@
Then, instead of using a fixed sequence of gates, a variational ansatz is used to parameterize the quantum layer, with the parameters being optimized during the training process. We used strongly entangling circuits from Pennylane [22], which apply rotations around the Z, Y, and Z axes, followed by entangling CNOT gates [23]. The following gates are defined by:
@@
For a single-qubit operation parameterized via Z-Y-Z rotations, the learnable unitary matrix is given by [10],[16]:
@@ III State-of-the-art
Quantum Machine Learning is a relatively new field in the Internet of Things (IoT) domain, and its use in aeronautics is even more recent. Work from [24] highlighted the use of unsupervised deep generative models by implementing two variational autoencoders (VAE) models for anomaly detection in commercial flight operations: one utilizing a factorized Bernoulli prior and another employing a restricted Boltzmann machine (RBM). This study proposes an RBM model compatible with a quantum annealer or Quantum Circuit Born Machine (QCBM). The results demonstrate a mean precision of 0.591 for the RBM model, displaying competitive performance when compared to its Gaussian counterpart, but the performance needs further improvement to achieve reliability. While the technique used in this research differs from the current study, it demonstrates a progressive approach to applying QC to aeronautic data. In the paper by [25], the authors present a deep learning method to evaluate the health state of the IMA. The QNN model is fed by extracted features from denoising autoencoders and classifies the data. Their method has been shown to be more effective and robust than the other four conventional algorithms, with an accuracy of 84.96% in the testing data.
One notable work on the use of hybrid models for network security is proposed in [8], in which the authors develop three hybrid quantum-classical frameworks based on quantum autoencoders. These models aim to recognize network anomalies using quantum deep learning, demonstrating strong performance across standard datasets with NISQ-era IBM quantum simulators. According to their results, all three frameworks display promising detection performance. Notably, the integration of QAEs with Quantum K-Nearest Neighbor (QkNN) was shown to have the highest accuracy among the three. Moreover, the work from [5] addresses the growing security concerns of ADS-B systems by proposing an anomaly detection system based on ML and DL techniques. In terms of results for their DL models, the multilayer perceptron achieved a competitive accuracy of 0.9817. In contrast to our work, which explores the potential of QML to handle anomaly detection in ADS-B data, the approach presented in [5] mainly focuses on establishing a public ADS-B dataset and enhancing the performance of classical ML on that dataset.
@@ IV-A Model #1: Hybrid-Fully Connected Quantum Neural Network (H-FQNN)
The H-FQNN model comprises four layers: three classical dense layers and one quantum layer. The first layer represents the input layer, which contains neurons equal to the number of features. The second layer is a hidden layer, also with neurons equal to the number of qubits, with a Rectified Linear Unit (ReLU) activation function. The third layer represents the quantum layer, also represented as a VQC. This quantum circuit contains qubits equal to the number of features, which are embedded with an R_X gate, passed through strongly entangling circuits, and measured using Pauli Z operators. However, we also varied the number of qubits in this layer to explore the outcome on the performance. Then, the final layer is the output layer, which contains 2 neurons. As BCEWithLogitsLoss and CrossEntropyLoss include respectively a sigmoid and a softmax layer internally, the output layer does not have an activation function as it would result in redundant computation. Figure 3 visually depicts the architecture of the H-FQNN. It is read from left to right, beginning with the classical input features, which are first processed by the classical layers and then passed to the quantum layer. The output of the quantum layer is then fed into the final dense layer for prediction, and the model is trained using the Adam optimizer [26].
@@ IV-C DATASET
In aeronautic research, the availability and diversity of commercial aircraft data are challenges that can slow down novel advancements [27]. To mitigate this difficulty and to focus on the impact of quantum properties, this research explored the Dataset 3 obtained from [5], which provided high-dimensional aircraft surveillance data for training and evaluation purposes. The dataset originates from the OpenSky Network, an open-access platform that provides ADS-B flight records [28]. The dataset comprises essential flight parameters related to aircraft positioning and flight dynamics, including latitude, longitude, barometric altitude, geometric altitude, velocity, and heading. It also includes aircraft identifiers in the ICAO24 format and a UNIX-formatted timestamps time. Dataset 3 was chosen among the available options because it contains basic trajectory manipulation attacks, including gradual drift in latitude/longitude as well as merge attacks that simulate complex flight patterns. Although these scenarios do not represent the full range of known ADS-B operational threats, the dataset provides an interpretable foundation for evaluating and isolating the contribution of QML methods in the study.
@@ IV-D Feature Selection
Following the feature selection process, we examined pairwise relationships between flight parameters to identify and eliminate highly correlated features while preserving those most relevant for anomaly detection. The analysis process can be visualized through a complete correlation matrix represented in Figure 5. In this matrix, lighter colors indicate stronger linear correlations between features. The diagonal elements of the matrix represent the self-correlation of each feature, which is always equal to 1. The other elements show the correlation between pairs of features. From this full feature representation, a threshold was established at 90% for parameter retention, eliminating strongly correlated variables while maintaining data integrity. We did not lower the threshold below 90% to ensure the retention of the velocity feature and at least one altitude feature, which are critical for flight dynamics and aircraft positioning [5]. Notably, the features barometric altitude and geoaltitude displayed perfect collinearity. barometric altitude was then removed, retaining only the latter as our altitude reference. This filtering process effectively reduced the dataset to six essential features.
@@ IV-E Data Preprocessing
To prepare the ADS-B dataset for training and simulation, we applied a structured preprocessing pipeline, inspired from the work done in [29][8], composed of sampling, normalization, dimensionality reduction, and encoding to ensure compatibility with both classical and QML models. First, to maintain class balance and control dataset size, we performed stratified sampling by selecting a fixed number of attack and normal flight samples based on different ratios. This ensured representative subsets for training, validating and testing while using imbalanced dataset. The sampled dataset is represented as:
@@ IV-F Evaluation Matrix
The anomaly detection for ADS-B data was evaluated using standard classification metrics, which compares the predicted classifications to the actual labels. The confusion matrix is a 2x2 table, shown in Figure 6, that summarizes the performance of a classification algorithm [30].

### v2

@@ I Introduction
Avionic systems have evolved significantly over the decades, transitioning from early federated architectures to today’s advanced Integrated Modular Avionics (IMA) frameworks. In the original federated model, each aircraft function—such as navigation, communication, or flight control—was managed by independent hardware and software components, leading to increased weight, complexity, and maintenance challenges. Modern IMA architectures have revolutionized this landscape by centralizing processing power through shared computing resources and modular software platforms, greatly enhancing flexibility, scalability, and fault tolerance [watkins2007ima]. This shift not only improves efficiency but also supports advanced communication protocols like Automatic Dependent Surveillance–Broadcast (ADS-B).
@@
In recent years, the adoption of ADS-B systems has led the aviation industry to an increase in the volume of aircraft surveillance data. The ADS-B system provides real-time data by transmitting the aircraft’ position, velocity and other flight characteristics to nearby aircraft and ground stations through a continuous, unencrypted data stream [varga2015adsb]. By leveraging satellite navigation, the ADS-B acquires precise positional information of an aircraft and is designed to improve aviation safety by enhancing situational awareness for pilots and air traffic controllers, aiding in collision avoidance and preventing runway incursions [manesh2017adsb]. Furthermore, it helps optimize routing, improving the overall efficiency of air traffic management. However, since aviation systems receive broadcast ADS-B data without validating the source or integrity, the system is vulnerable to cyber threats such as spoofing, ghost aircraft injection, trajectory manipulation, and other attacks through the transmission of false data into air traffic control systems without validation [manesh2017adsb],[kim2017adsb].
@@
Traditionally, a range of classical Machine Learning (ML) techniques have been employed in ADS-B security to detect anomalies and improve data validation. Algorithms such as Decision Trees, Support Vector Machines (SVM) and Neural Networks (NN) have demonstrated success in identifying and mitigating cyber threats in ADS-B data stream [Cevi2024Anomaly],[Pirolley2024GhostAircraft]. However, these classical methods often involve high computational complexity and memory usage and may lose critical temporal information during feature transformation. These limitations highlight the need for more adaptive and scalable approaches to secure ADS-B communication, especially as air traffic surveillance data grows in complexity and volume.
@@
In parallel, as quantum hardware continues to improve, the growing field of QML has shown promising potential in accelerating the processing speed and effectively handling the high dimensionality associated with complex datasets through the quantum properties of superposition and entanglement [Torabian2023Compositional],[Hdaib2024]. Hybrid quantum-classical models, which combine quantum circuits with classical neural networks, offer the potential to enhance learning capabilities while remaining compatible with near-term noisy intermediate-scale quantum (NISQ) devices.
@@
In this work, we explore the feasibility and application of a hybrid quantum-classical neural network, a H-FQNN, to perform anomaly detection on trajectory-based spoofing ADS-B data. We construct two quantum circuit-based hybrid models using one-hot encoded labels and binary cross-entropy loss and compare both to their classical counterpart, an FNN. The models are trained and evaluated on a publicly available ADS-B anomaly dataset that includes gradual manipulations of latitude and longitude, as well as merge attacks simulating new and complex flight patterns [Cevi2024Anomaly]. Detecting such trajectory anomalies is critical in operational aviation environments, as they may lead to false traffic alerts and in-flight disruptions, potentially resulting in serious safety consequences [Realities_adsb_martin].
@@ II-A Quantum Computing (QC)
Quantum computing can be understood as a computational paradigm that leverages the principles of quantum mechanics and is built upon the fundamental unit of quantum information, the qubit [nielsen2010quantum]. Often represented as the spin of an electron or the polarization of a photon [bernhardt2019quantum] [hennig2024introducing], the qubit is the unit of computation with the basis states of:
@@
The fundamental utility of the quantum systems lies in the ability to represent information using the quantum mechanical property of entanglement and superposition. Superposition is the property that enables a quantum system to exist in multiple states simultaneously until measured. This can be represented as the linear combinations of states, written $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$ where $\alpha$ and $\beta$ are complex coefficients for which the square of their magnitudes, $|\alpha|^{2}$ and $|\beta|^{2}$, give the probabilities of measuring the system in the $|0\rangle$ and $|1\rangle$ states. Entanglement is another elementary property defining a group of qubits which are interconnected and cannot be described independently. In a system of 2 entangled qubits, it can be interpreted as the state of one qubit directly influencing the state of the other, with the famous example of the Bell state [bell1964einstein] [qipc2021]. A simple representation of the equation is:
@@ II-B Quantum Machine Learning (QML)
The convergence of Machine Learning and Quantum Computing has the potential to improve the communication technologies of the space and aeronautical sectors[Brin2024]. Quantum Machine Learning brings together the current state of the art in Machine Learning and harnesses the quantum mechanical properties of entanglement and superposition to attempt to create better performing models. There are four different approaches to combine QC and ML [Schuld2021] [joshi2021evaluating] as illustrated in Figure 1. Classical input-Classical model (CC) refers to standard classical machine learning, where both data and models are classical. Classical input-Quantum model (CQ) involves processing classical data using quantum models, such as parameterized quantum circuits. Quantum input-Classical model (QC) describes the use of classical machine learning techniques to analyze data generated by quantum systems. Finally, Quantum input-Quantum model (QQ) leverages both quantum data and quantum models, enabling fully quantum learnable frameworks. These four paradigms define the various ways in which quantum and classical computing can synergize to advance machine learning capabilities. In this article, we refer to a hybrid model as a CQ framework that implements a quantum algorithm using classical input data.
Figure 1: Four different approaches to combine the disciplines of quantum computing and machine learning. (The figure is adapted from Schuld and Petruccione [Schuld2021])
@@ II-C Variational Quantum Classifier (VQC)
A Variational Quantum Classifier can be considered as the quantum equivalent of a layer in a hybrid NN model. VQCs are particularly well suited for NISQ computers as the process of optimizing the cost function of a machine learning algorithm also helps mitigate the inherent decoherence and errors present in current quantum hardware [Havlicek2019]. The quantum circuit of a VQC begins with an initial state of qubits and uses a parameterized quantum circuit, known as a feature map, to encode features as quantum information [Coles2021] [Kruse2023Variational]. In this article, we used the angle embedding method, which encodes classical information of n features into the rotation angle of n qubits, effectively manipulating information from the vector space $\mathbb{C}^{N}$, where $N=2^{n}$ is the dimension of the Hilbert space corresponding to n qubits [Lloyd2020]. Specifically, we used $R_{X}$ gates, which apply a rotation of $\theta$ around the x-axis of a qubit, mapping the classical feature $x\in\mathbb{R}$ to the quantum state $|\psi(x)\rangle\in\mathcal{H}$, where $\mathcal{H}$ is a two-dimensional Hilbert space. The $R_{X}$ gate is defined by:
@@
Then, instead of using a fixed sequence of gates, a variational ansatz is used to parameterize the quantum layer, with the parameters being optimized during the training process. We used strongly entangling circuits from Pennylane [pennylane_strongly_entangling_layers], which apply rotations around the Z, Y, and Z axes, followed by entangling CNOT gates [schuld2018circuit]. The following gates are defined by:
@@
For a single-qubit operation parameterized via Z-Y-Z rotations, the learnable unitary matrix is given by [nielsen2010quantum],[Schuld2021]:
@@ III State-of-the-art
Quantum Machine Learning is a relatively new field in the Internet of Things (IoT) domain, and its use in aeronautics is even more recent. Work from [Templin2023Anomaly] highlighted the use of unsupervised deep generative models by implementing two variational autoencoders (VAE) models for anomaly detection in commercial flight operations: one utilizing a factorized Bernoulli prior and another employing a restricted Boltzmann machine (RBM). This study proposes an RBM model compatible with a quantum annealer or Quantum Circuit Born Machine (QCBM). The results demonstrate a mean precision of 0.591 for the RBM model, displaying competitive performance when compared to its Gaussian counterpart, but the performance needs further improvement to achieve reliability. While the technique used in this research differs from the current study, it demonstrates a progressive approach to applying QC to aeronautic data. In the paper by [Gao2018IMA], the authors present a deep learning method to evaluate the health state of the IMA. The QNN model is fed by extracted features from denoising autoencoders and classifies the data. Their method has been shown to be more effective and robust than the other four conventional algorithms, with an accuracy of 84.96% in the testing data.
One notable work on the use of hybrid models for network security is proposed in [Hdaib2024], in which the authors develop three hybrid quantum-classical frameworks based on quantum autoencoders. These models aim to recognize network anomalies using quantum deep learning, demonstrating strong performance across standard datasets with NISQ-era IBM quantum simulators. According to their results, all three frameworks display promising detection performance. Notably, the integration of QAEs with Quantum K-Nearest Neighbor (QkNN) was shown to have the highest accuracy among the three. Moreover, the work from [Cevi2024Anomaly] addresses the growing security concerns of ADS-B systems by proposing an anomaly detection system based on ML and DL techniques. In terms of results for their DL models, the multilayer perceptron achieved a competitive accuracy of 0.9817. In contrast to our work, which explores the potential of QML to handle anomaly detection in ADS-B data, the approach presented in [Cevi2024Anomaly] mainly focuses on establishing a public ADS-B dataset and enhancing the performance of classical ML on that dataset.
@@ IV-A Model #1: Hybrid-Fully Connected Quantum Neural Network (H-FQNN)
The H-FQNN model comprises four layers: three classical dense layers and one quantum layer. The first layer represents the input layer, which contains neurons equal to the number of features. The second layer is a hidden layer, also with neurons equal to the number of qubits, with a Rectified Linear Unit (ReLU) activation function. The third layer represents the quantum layer, also represented as a VQC. This quantum circuit contains qubits equal to the number of features, which are embedded with an R_X gate, passed through strongly entangling circuits, and measured using Pauli Z operators. However, we also varied the number of qubits in this layer to explore the outcome on the performance. Then, the final layer is the output layer, which contains 2 neurons. As BCEWithLogitsLoss and CrossEntropyLoss include respectively a sigmoid and a softmax layer internally, the output layer does not have an activation function as it would result in redundant computation. Figure 3 visually depicts the architecture of the H-FQNN. It is read from left to right, beginning with the classical input features, which are first processed by the classical layers and then passed to the quantum layer. The output of the quantum layer is then fed into the final dense layer for prediction, and the model is trained using the Adam optimizer [kingma2017adammethodstochasticoptimization].
@@ IV-C DATASET
In aeronautic research, the availability and diversity of commercial aircraft data are challenges that can slow down novel advancements [Li2019Reviewing]. To mitigate this difficulty and to focus on the impact of quantum properties, this research explored the Dataset 3 obtained from [Cevi2024Anomaly], which provided high-dimensional aircraft surveillance data for training and evaluation purposes. The dataset originates from the OpenSky Network, an open-access platform that provides ADS-B flight records [Schaefer2014]. The dataset comprises essential flight parameters related to aircraft positioning and flight dynamics, including latitude, longitude, barometric altitude, geometric altitude, velocity, and heading. It also includes aircraft identifiers in the ICAO24 format and a UNIX-formatted timestamps time. Dataset 3 was chosen among the available options because it contains basic trajectory manipulation attacks, including gradual drift in latitude/longitude as well as merge attacks that simulate complex flight patterns. Although these scenarios do not represent the full range of known ADS-B operational threats, the dataset provides an interpretable foundation for evaluating and isolating the contribution of QML methods in the study.
@@ IV-D Feature Selection
Following the feature selection process, we examined pairwise relationships between flight parameters to identify and eliminate highly correlated features while preserving those most relevant for anomaly detection. The analysis process can be visualized through a complete correlation matrix represented in Figure 5. In this matrix, lighter colors indicate stronger linear correlations between features. The diagonal elements of the matrix represent the self-correlation of each feature, which is always equal to 1. The other elements show the correlation between pairs of features. From this full feature representation, a threshold was established at 90% for parameter retention, eliminating strongly correlated variables while maintaining data integrity. We did not lower the threshold below 90% to ensure the retention of the velocity feature and at least one altitude feature, which are critical for flight dynamics and aircraft positioning [Cevi2024Anomaly]. Notably, the features barometric altitude and geoaltitude displayed perfect collinearity. barometric altitude was then removed, retaining only the latter as our altitude reference. This filtering process effectively reduced the dataset to six essential features.
@@ IV-E Data Preprocessing
To prepare the ADS-B dataset for training and simulation, we applied a structured preprocessing pipeline, inspired from the work done in [Payares2021Quantum][Hdaib2024], composed of sampling, normalization, dimensionality reduction, and encoding to ensure compatibility with both classical and QML models. First, to maintain class balance and control dataset size, we performed stratified sampling by selecting a fixed number of attack and normal flight samples based on different ratios. This ensured representative subsets for training, validating and testing while using imbalanced dataset. The sampled dataset is represented as:
@@ IV-F Evaluation Matrix
The anomaly detection for ADS-B data was evaluated using standard classification metrics, which compares the predicted classifications to the actual labels. The confusion matrix is a 2x2 table, shown in Figure 6, that summarizes the performance of a classification algorithm [Ting2010].